Circular No. 6260

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

AX J2315–592 K. Misaki, Y. Kamata, Y. Terashima, K. Isobe, H. Kunieda, and Y.

Tawara, Nagoya University; T. Tsuru, Y. Maeda, M. Sakano, M. Akiyama, and H. Sogawa, Kyoto University; and M. Ishida and R. Fujimoto, Institute of Space and Astronautical Science, communicate: “ASCA discovered a new x-ray source during the observation of IRAS 23128–5919 carried out during Nov. 2.55–3.87 UT. The new source is located at α = 23^{h}15^{m}18^{s}, δ = −59^{o}10′.7 (equinox 2000.0; error circle 1′ radius). There are two optical stars (at 500 nm) of mag 16–17 within this circle in the online chart of ‘Sky View’ (provided by NASA). Unfortunately, the source was out of the field-of-view of the SIS, and only the GIS data are available. The average x-ray flux is 2.3 × 10^{−11} erg cm^{−2} s^{−1}, or 1.1 mCrab, at 2–10 keV. No xray source with this intensity is known in the HEAO-1 catalogue, GINGA archival data, ROSAT WFC catalogue, or SIMBAD. The new source AX J2315–592 shows a strong periodic x-ray modulation with period 5360 ± 100 s (or 1.49 ± 0.03 hr). The lightcurve is characterized by a deep and narrow intensity minimum, which lasts for 20 percent of the period. The deficit of the countrate is ≈ 90 and 70 percent of the peak intensity in the energy ranges 0.7–2.4 and 2.4–10 keV, respectively. The average spectrum is hard and is represented either by a power law with photon index between 1.3 and 1.4 or by thermal bremsstrahlung with a temperature of > 20 keV. The absorption is small and the line-of-sight hydrogen column density is < 1 × 10^{21} cm^{−2}. A prominent iron emission line is detected at 6.79 ± 0.06 keV with an equivalent width of ∼ 1000 eV. The line is broad with 1σ width of 200–400 eV, suggestive of a blend of thermal and fluorescent emissions. From the lightcurve and spectral characteristics, AX J2315–592 is probably a new AM Her-type cataclysmic variable. Follow-up optical observations (spectroscopy and polarimetry) are encouraged.”

R CORONAE BOREALIS Visual magnitude estimates (cf. IAUC 6245): Oct. 19.14 UT, 10.6 (C.

Spratt, Victoria, BC); 20.72, 11.5 (L. Szentaskó, Veresegyhaz, Hungary); 23.783, 11.8 (P. Schmeer, Bischmisheim, Germany); 26.74, 11.9 (M. V. Zanotta, Laino, Italy); 31.74, 12.9 (Zanotta); Nov. 2.75, 12.1 (Szentaskó); 5.74, 13.3 (Zanotta); 9.72, 13.2 (A. Mizser, Budapest, Hungary).

1995 November 11 Daniel W. E. Green